# Mitigating the Alignment Tax of RLHF

### v3

@@ 4 Investigating the Mysteries of Model Averaging for Alleviating Alignment Tax
$\small\xi^{(1)}-\xi^{(2)}$ $=F_{p}\left(\frac{\sqrt{2}(1-p)n}{\sqrt{n+n_{o}}}\right)-F_{p}\left((1-p)\sqrt{n}\right)\geq 0,$
where the equality holds when $n_{o}=n$. Here $F_{p}(x)$ is a cumulative density function increasing with $x$.
@@
Proposition 4.1 demonstrates that when $\mathcal{T}_{a}$ and $\mathcal{T}_{b}$ are more similar, the averaging of models ($f_{a}$ and $f_{b}$) yields greater improvement. However, this improvement is reduced if $f_{a}$ and $f_{b}$ use more overlapping features. Recall that each weak feature can fail with probability $p$. If $\mathcal{T}_{a}$ and $\mathcal{T}_{b}$ are similar, the feature utilized by the two models would be projected into a shared space, allowing model averaging to take advantage of a more diverse set of features. This diversity reduces the probability of model failure because a diverse set of features is less likely to fail together simultaneously (Lin et al., 2023). However, if $\mathcal{T}_{a}$ and $\mathcal{T}_{b}$ are dissimilar, for example, if $|\mathcal{Y}_{a}\cap\mathcal{Y}_{b}|=0$ and the feature spaces corresponding to $\mathcal{Y}_{a}$ and $\mathcal{Y}_{b}$ are disjoint, then the features in the space of $\mathcal{Y}_{a}$ would not provide any information for predicting $\mathcal{Y}_{b}$. Therefore, averaging $f_{a}$ and $f_{b}$ would not improve the prediction of either task in this case. Refer to Appendix E.3 for a detailed discussion.
@@
The results of the above scheme are denoted as “Input Part MA”. “Middle Part MA” and “Output Part MA” represent that we average the middle and output parts, respectively. Figure 7 illustrates that the reward-tax trade-off varies distinctly when different parts of the transformers are averaged. Specifically, when we average the low-level layers, we observe a “magical” improvement in both the NLP tasks and alignment rewards, which is consistent with our previous analysis. Furthermore, we show results in Appendix D.2 that the magical improvement in averaging the low-level parts is consistent among DPO and PPO models.
5 Adaptive Model Averaging
@@
Figure 6: Illustration of proof of concept experiments. We divide the Transformer into 3 parts. We only average one part each time.
Figure 6: Illustration of proof of concept experiments. We divide the Transformer into 3 parts. We only average one part each time.
Figure 7: Merging different parts of the transformers.
@@
We have already shown that averaging different layers results in diverse patterns of reward-tax trade-off (Wu et al., 2022; Lee et al., 2022b). Therefore, different layers should not be equally treated during averaging. This leads to a natural question: can we enhance the reward-tax trade-off by using adaptive weights for different layers? Consequently, we conduct proof-of-concept experiments to provide affirmative answers to this question and subsequently propose a practical algorithm.
@@
Proof of Concept. The following proof of concept experiments provide insights into average different layers with various ratios. We use different averaging ratio, i.e., $\alpha_{1},\alpha_{2},\alpha_{3}$, for the three parts. Specifically, the $i$th part of the averaged model is simply $\alpha_{i}\theta^{[i]}+(1-\alpha_{i})\theta_{0}^{[i]}$. We try three patterns experiment given a base $\alpha\in\{0.2,0.3,0.4\}$ : (a) $\alpha_{1}=\alpha_{2}=\alpha_{3}=\alpha$; (b) $\alpha_{1}=\alpha_{2}=\alpha$, $\alpha_{3}=\alpha-0.1$, and (c) $\alpha_{1}=\alpha$, $\alpha_{2}=\alpha_{3}=\alpha-0.1$. We use $(\alpha|\alpha|\alpha)$, $(\alpha|\alpha|\alpha-0.1)$ and $(\alpha|\alpha-0.1|\alpha-0.1)$ to denote these three patterns, respectively. These results confirm that certain ratio combinations exceed the trade-off curve of vanilla model averaging, as displayed in Figure 8. Notably, some combination ratios consistently outperform the equal ratio across various benchmarks. This affirms the potential to identify consistent combination ratios that demonstrate superior performance across a broad spectrum of benchmarks in terms of reward-tex trade-off.
Figure 8: Evaluation of different combination ratios.
@@
Adaptive Model Averaging. Upon dividing the Transformer into $K$ parts, our objective is to adaptively determine a combination ratio for different layers that consistently perform well across an extensive range of tasks. The conventional averaging method uses a shared $\alpha$ for all layers, playing a crucial role in defining the trade-off between reward and tax. We aim to identify an optimized combination of $(\alpha_{1},...,\alpha_{K})$ to replace a uniform $\alpha$. Let $\theta(K)$ represent the model merged by $(\alpha_{1},...,\alpha_{K})$. In particular, the $k$th component of the merged model $\theta(K)$ is given by
$\theta^{[k]}(K):=\alpha_{k}\theta^{[k]}+(1-\alpha_{k})\theta_{0}^{[k]},\forall k\in 1,...,K.$ (5)
@@
The intuition behind AMA is outlined as follows:
  • When maintaining the mean, i.e., $\frac{1}{K}\sum_{k}\alpha_{k}$, as $\alpha$, we can compare AMA performance with the performance of vanilla model averaging with the same $\alpha$.
@@
  • We only optimize $K$ parameters, where $K$ is typically small. For example, we adopt $K=3$ by default and also include results with varying $K$ to the ablation study. This helps to ensure that the forgetting level of $(\alpha_{1},...,\alpha_{K})$ remains close to $\alpha$. Intuitively, if we optimize a large number of parameters, it could easily lead to over-fitting in the in-domain (RLHF reward) and may also result in more significant forgetting.
The whole algorithm is summarized as follows:
Algorithm 1 AMA: Adaptive Model Averaging
0:  The output policy $\pi_{\theta(K)}$.
1:  Perform vanilla RLHF by Eqn (1) and obtain $\pi_{\theta}$.
2:  Distill $\mathcal{D}_{\theta}$ from $\pi_{\theta}$ according to Eqn. (7).
4:  Obtain the averaged model $\theta(K)$ with $\alpha_{1},...,\alpha_{K}$ according to Eqn. (5).
5:  Solve adaptive ratios $\alpha_{1},...,\alpha_{K}$ according to Eqn. (8).
6:  Return the $\theta(K)$ with the optimized $\alpha_{1},...,\alpha_{K}$.
@@
Results. The results of AMA are shown in Figure 9. We can see that AMA can consistently push forward the Perato-front of the vanilla model averaging. Furthermore, such improvement is consistent over various RLHF algorithms. More detailed results (e.g., on Commonsense QA and Translation with different RLHF algorithms) of AMA can be found in Appendix D.4.
Figure 9: Results of Adaptive Model Averaging (AMA). (Left) RSF, (Middle) DPO, (right) PPO. We show $\alpha\in[0.1,0.6]$ for better visualization. Refer to Appendix D.4 for the full results.
Ablation results on different $K$. We tested different values of $K$ with $\alpha=0.2,0.4,0.6$ as illustrated in Figure 11. The trade-off curve shows a slight decrease as we increase $K$ from 3 to 6 and 9, but still consistently improves over the vanilla model averaging. This decrease is likely due to overfitting. Specifically, comparing the performance of AMA with different $K$ for the same mean ratio, we observe that as the alignment reward increases with an increase in $K$ from 3 to 9, the reading comprehension performance drops. This suggests that increasing $K$ causes AMA to over-fit the alignment reward, leading to increased alignment tax.
Further verification on Mistral-7B based models. To further validate our method on larger LLMs, e.g., Mistral-7B (Jiang et al., 2023a) based models, we apply model averaging (MA) and adaptive model averaging (AMA) on Zephyr-7B-$\beta$ [^3] https://huggingface.co/HuggingFaceH4/zephyr-7b-beta (Tunstall et al., 2023) which is trained with DPO on the SFT version, Mistral-7B-SFT-$\beta$ [^4] https://huggingface.co/HuggingFaceH4/mistral-7b-sft-beta.
@@
Figure 10: Results of AMA with different choices of $K$.
Figure 10: Results of AMA with different choices of $K$.
Figure 11: Results of Zephyr-7B-$\beta$
@@
Here we use the PairRM (Jiang et al., 2023b) to judge the helpfulness and evaluate models on AlpacaEval 2.0 (Li et al., 2023). We take the win rates to measure the helpfulness of different language models. The results range across $\alpha=0,0.2,\dots,1.0$ depicted in Figure 11 demonstrate that MA effectively achieves a strong Pareto front to mitigate forgetting in the Mistral-7B models. Additionally, our AMA algorithm shows further improvement compared to the MA method. These findings provide additional evidence of the generalization ability of our AMA model.
@@ 6 Conclusion
In this paper, we conducted a systematic comparison of existing methods for alleviating alignment tax and found that model averaging achieves a notably strongest Pareto front. The success of model averaging could be attributed to its ability to enhance feature diversity, as shown in the theoretical analysis. Additionally, we demonstrate that different model parts should not be equally weighted and propose the Adaptive Model Averaging (AMA) framework. Our extensive experiments confirm the effectiveness of the AMA framework.
Acknowledgements
We are grateful for the valuable insights and engaging discussions with Gangwei Jiang, Ying Wei, Mitchell Wortsman, and Ludwig Schmidt.
@@ Appendix B RLHF Basics
PPO is the the classical method for RLHF and has gained its success in aligning Chat-GPT (OpenAI, 2023). In contrast to the implementation in traditional DRL scenario, for alignment of LLMs, following (Ziegler et al., 2019; Wu et al., 2021; Ouyang et al., 2022; Rafailov et al., 2023; Liu et al., 2023), we modify the reward optimization as the following KL-regularized form:
@@
Appendix C Implementation Details

### v4

@@ 5 Unravelling the Mysteries of Model Averaging for Alleviating Alignment Tax
$\xi^{(1)}-\xi^{(2)}=$ $F_{p}\left(\frac{\sqrt{2}(1-p)n}{\sqrt{n+n_{o}}}\right)$
$-F_{p}\left((1-p)\sqrt{n}\right)\geq 0,$
where the equality holds when $n_{o}=n$ and $F_{p}(x)$ is a cumulative density function in Appendix F.4.
@@
Proposition 5.1 demonstrates that when $\mathcal{T}_{a}$ and $\mathcal{T}_{b}$ are more similar, the averaging of models ($f_{a}$ and $f_{b}$) yields greater improvement. However, this improvement is reduced if $f_{a}$ and $f_{b}$ use more overlapping features. Recall that each weak feature can fail with probability $p$. If $\mathcal{T}_{a}$ and $\mathcal{T}_{b}$ are similar, the feature utilized by the two models would be projected into a shared space, allowing model averaging to take advantage of a more diverse set of features. This diversity reduces the probability of model failure because a diverse set of features is less likely to fail together simultaneously (Lin et al., 2023). However, if $\mathcal{T}_{a}$ and $\mathcal{T}_{b}$ are dissimilar, for example, if $|\mathcal{Y}_{a}\cap\mathcal{Y}_{b}|=0$ and the feature spaces corresponding to $\mathcal{Y}_{a}$ and $\mathcal{Y}_{b}$ are disjoint, then the features in the space of $\mathcal{Y}_{a}$ would not provide any information for predicting $\mathcal{Y}_{b}$. Therefore, averaging $f_{a}$ and $f_{b}$ would not improve the prediction of either task in this case. Refer to Appendix F.3 for a detailed discussion.
@@
The results of the above scheme are denoted as “Input Part MA". “Middle Part MA" and “Output Part MA" represent that we average the middle and output parts, respectively. Figure 4 illustrates that the alignment-forgetting trade-off varies distinctly when different parts of the transformers are averaged. Specifically, when we average the low-level layers, we observe a “magical” improvement in both the NLP tasks and alignment rewards, which is consistent with our previous analysis. Furthermore, we show results in Appendix E.2 that the magical improvement in averaging the low-level parts is consistent among DPO and PPO models.
6 Heterogeneous Model Averaging
@@
Figure 4: (Left) Illustration of proof of concept experiments. We divide the Transformer into 3 parts. We only average one part each time. (Right) Merging different parts of the transformers.
@@
We have already shown that averaging different layers results in diverse patterns of alignment-forgetting trade-off (Wu et al., 2022; Lee et al., 2022b). Therefore, different layers should not be equally treated during averaging. This leads to a natural question: can we enhance the alignment-forgetting trade-off by using adaptive weights for different layers? Consequently, we conduct proof-of-concept experiments to provide affirmative answers to this question and subsequently propose a practical algorithm.
@@
Proof of Concept. The following proof of concept experiments provide insights into average different layers with various ratios. We use different averaging ratio, i.e., $\alpha_{1},\alpha_{2},\alpha_{3}$, for the three parts. Specifically, the $i$th part of the averaged model is simply $\alpha_{i}\theta^{[i]}+(1-\alpha_{i})\theta_{0}^{[i]}$. We try three patterns experiment given a base $\alpha\in\{0.2,0.3,0.4\}$ : (a) $\alpha_{1}=\alpha_{2}=\alpha_{3}=\alpha$; (b) $\alpha_{1}=\alpha_{2}=\alpha$, $\alpha_{3}=\alpha-0.1$, and (c) $\alpha_{1}=\alpha$, $\alpha_{2}=\alpha_{3}=\alpha-0.1$. We use $(\alpha|\alpha|\alpha)$, $(\alpha|\alpha|\alpha-0.1)$ and $(\alpha|\alpha-0.1|\alpha-0.1)$ to denote these three patterns, respectively. These results confirm that certain ratio combinations exceed the trade-off curve of vanilla model averaging, as displayed in Figure 9 in Appendix C.3. Notably, some combination ratios consistently outperform the equal ratio across various benchmarks. This affirms the potential to identify consistent combination ratios that demonstrate superior performance across a broad spectrum of benchmarks in terms of alignment-forgetting trade-off.
@@
Heterogeneous Model Averaging. Upon dividing the Transformer into $K$ parts, our objective is to adaptively determine a combination ratio for different layers that consistently perform well across an extensive range of tasks. The conventional averaging method uses a shared $\alpha$ for all layers, playing a crucial role in defining the trade-off between reward and tax. We aim to identify an optimized combination of $(\alpha_{1},...,\alpha_{K})$ to replace a uniform $\alpha$. Let $\theta(K)$ represent the model merged by $(\alpha_{1},...,\alpha_{K})$. In particular, the $k$th component of the merged model $\theta(K)$ is given by
$\theta^{[k]}(K):=\alpha_{k}\theta^{[k]}+(1-\alpha_{k})\theta_{0}^{[k]},\forall k\in 1,...,K.$
@@
The intuition behind HMA is outlined as follows: (1) When maintaining the mean, i.e., $\frac{1}{K}\sum_{k}\alpha_{k}$, as $\alpha$, we can compare HMA performance with the performance of vanilla model averaging with the same $\alpha$. (b) We only optimize $K$ parameters, where $K$ is typically small. For example, we adopt $K=3$ by default and also include results with varying $K$ to the ablation study. This helps to ensure that the forgetting level of $(\alpha_{1},...,\alpha_{K})$ remains close to $\alpha$. Intuitively, if we optimize a large number of parameters, it could easily lead to over-fitting in the in-domain (RLHF reward) and may also result in more significant forgetting. The whole algorithm is summarized Algorithm 1 in appendix.
@@
Results. The results of HMA are shown in Figure 5. We can see that HMA can consistently push forward the Perato-front of the vanilla model averaging. Furthermore, such improvement is consistent over various RLHF algorithms. More detailed results (e.g., on Commonsense QA and Translation with different RLHF algorithms) of HMA can be found in Appendix E.5.
@@
Figure 5: Results of our HMA. (Top) HMA for RSF ( $\alpha\in[0.1,0.6]$), (Bottom) HMA for DPO ( $\alpha\in[0.1,0.6]$). (Right) HMA for RSF with different choices of $K$. Refer to Appendix E.5 for more results.
Ablation results on different $K$. We tested different values of $K$ with $\alpha=0.2,0.4,0.6$ as illustrated in Figure 5 (Right). The trade-off curve shows a slight decrease as we increase $K$ from 3 to 6 and 9, but still consistently improves over the vanilla model averaging. This decrease is likely due to overfitting. Specifically, comparing the performance of HMA with different $K$ for the same mean ratio, we observe that as the alignment reward increases with an increase in $K$ from 3 to 9, the reading comprehension performance drops.
How to choose the averaging ratio. In practice, we determine the averaging ratio $\alpha$ for adopting vanilla MA or our HMA. Changing the averaging ratio for MA and HMA is convenient as these methods are applied after training the vanilla RLHF checkpoint. The comprehensive results in Figures 3, 5, and 16 (details in Appendix C.4) show that $\alpha=0.2$ can consistently alleviate the alignment tax without hurting alignment performance. Further results of Zephyr-7B are shown in Figure 6. Additionally, the performance of the averaging ratio on different benchmarks (Figure 9) exhibits similar trends. Hence, we believe $\alpha=0.2$ is a suitable choice that can generalize to more tasks.
Figure 6: Results of Zephyr-7B-$\beta$ evaluated by open sourced preference model. (Top) Similar trends evaluated by PairRM when we average different blocks. (Bottom) Our HMA consistently improve over MA.
Model Win-Rate Reading CommonSense Trans
Zephyr-7B-$\beta$ 8.10% 37.47 66.34 36.55
HMA (Ours) 9.32% 38.93 66.55 37.23
Zephyr-7B-Gemma 11.3% 41.15 66.3 38.09
HMA (Ours) 11.5% 42.45 66.4 38.71
Table 1: GPT4 evaluation of experiments of Zephyr-7B-$\beta$ and Zephyr-7B-Gemma on Alpaca benchmark. Reading is short for Reading Comprehension, which is evaluated by F1. CommonSence is evaluated by Accuracy (%). Trans is short for Translation Fr-En, evaluated by BLEU.
@@
Other models results. To further validate our method on larger LLMs, e.g., Mistral-7B (Jiang et al., 2023a) based models, we apply model averaging (MA) and Heterogeneous Model Averaging (HMA) on Zephyr-7B-$\beta$ [^5] https://huggingface.co/HuggingFaceH4/zephyr-7b-beta (Tunstall et al., 2023) which is trained with DPO on the SFT version, Mistral-7B-SFT-$\beta$ [^6] https://huggingface.co/HuggingFaceH4/mistral-7b-sft-beta. We also apply HMA on Zephyr-7B-Gemma  [^7] https://huggingface.co/HuggingFaceH4/zephyr-7b-gemma-v0.1 which is aligned based on Gemma-7B [^8] https://huggingface.co/google/gemma-7b model. Here we use the the publicly available preference model PairRM (Jiang et al., 2023b) to judge the helpfulness and evaluate models on AlpacaEval 2.0 (Li et al., 2023). We reports the win rates of each model. Figure 6 (Top) shows that the trends of averaging different layers evaluated by PairRM are similar with the results evaluated by our own reward model. The results range across $\alpha=0,0.2,\dots,1.0$ depicted in Figure 6 (Bottom) demonstrate that MA effectively achieves a strong Pareto front to mitigate forgetting in the Mistral-7B models. Additionally, our HMA algorithm shows further improvement compared to the MA method.
GPT4 Evaluation. We also use GPT4 to evaluate HMA on AlpacaEval 2.0 (Li et al., 2023). Due to the limited quota, we only compare HMA with $\alpha=0.2$ with vanilla Zephyr-7B-$\beta$ ($\alpha=0.2$ is recommended by the previous discussion). In Table 1, we summarize their Win-Rate against GPT4 as well as their performance on NLP tasks. We show that HMA consistently outperforms Zephyr-7B-$\beta$ on all the metrics.
@@ 7 Conclusion
In this paper, we highlight the surprisingly effectiveness of model averaging and propose the Heterogeneous Model Averaging (HMA) framework to further enhance the performance.
Limitations
Though our HMA significantly alleviates the alignment tax, it has not been fully eliminated. Future work could explore the theoretical lower bound of the alignment tax and determine which method could achieve the optimal trade-off.
@@ Appendix A Related Work
Alignment tax.
Ouyang et al. (2022) reports that they observe significant alignment tax when developing InstructGPT. They have also tried to adopt Experience Replay to alleviate this issue, which is followed by Zheng et al. (2023). However, we show in Appendix C.1 that Experience Relay is less favorable when compared with model averaging. Noukhovitch et al. (2024) tried to use stochastic weight averaging, which still under-performs our method as shown in Figure 3. Li et al. (2024) finds that DPO induces less alignment tax compared with other RLHF algorithms, which is consistent with our findings (e.g., Figure 5). Askell et al. (2021) reports that they didn’t observe significant alignment tax when prompting LLM to align with humans. However, we focus on a more standard setting that the LLM is fully fine-tuned for RLHF.
@@ Appendix B RLHF Basics
PPO is the the classical method for RLHF and has gained its success in aligning Chat-GPT (OpenAI, 2023). In contrast to the implementation in traditional DRL scenario, for alignment of LLMs, following (Ziegler et al., 2019; Wu et al., 2021a; Ouyang et al., 2022; Rafailov et al., 2023; Liu et al., 2023), we modify the reward optimization as the following KL-regularized form:
@@
B.1 Algorithm of Heterogeneous Model Averaging
The algorithm of Heterogeneous Model Averaging is summarized as follows:
Algorithm 1 HMA: Heterogeneous Model Averaging
0:  The output policy $\pi_{\theta(K)}$.
1:  Perform vanilla RLHF by Eqn (1) and obtain $\pi_{\theta}$.
2:  Distill $\mathcal{D}_{\theta}$ from $\pi_{\theta}$ according to Eqn. (8).
4:  Obtain the averaged model $\theta(K)$ with $\alpha_{1},...,\alpha_{K}$.
5:  Solve Heterogeneous ratios $\alpha_{1},...,\alpha_{K}$ according to Eqn. (9).
6:  Return the $\theta(K)$ with the optimized $\alpha_{1},...,\alpha_{K}$.
Appendix C More Results
C.1 Experience Replay
In our alignment tax situation, we aim to preserve a wide range of abilities gained during pre-training. It is possible to replay a small subset of pretraining data, which also known as Experience Replay (ER) (Rebuffi et al., ; Shin et al., 2017). However, this method is less practical since pre-training datasets of most models are often not publicly available. Further more, even if we can access the pre-training data, retaining a subset of the pre-training data entails extra computational costs and implementation intricacies, making it less preferable (Noukhovitch et al., 2023). In this part, we compare ER with MA. Specifically, we include a small proportion of randomly subsampled pre-training data during the RLHF stage. Here, we denote $\mathcal{D}_{{pre}}$ as the pre-training data distribution, and our objective is to solve the following:
We experiment with different penalty weights $\lambda$ such as $0.25$, $0.5$, $1$, $2$, and $4$. Importantly, we utilize the data proportion as a proxy for setting the penalty weight. For instance, we do not explicitly apply a penalty of 4 when $\lambda=4$; instead, we include 4 times the replay data over the RLHF data in a batch. Refer to the Appendix D for more details.
Figure 7: Comparison of model averaging with Experience Replay.
C.2 Reward Penalty
It is a common practice to impose Kullback–Leibler (KL) penalty on the RL reward in the PPO. Such a penalty can also regularize the policy to stay closer to the initial policy, which in return can reduce the alignment tax. Following (Ziegler et al., 2019; Wu et al., 2021a; Ouyang et al., 2022; Yuan et al., 2023), we modify the raw reward function with an additional KL penalty (Ziegler et al., 2019).
where we use $\mathrm{KL}(\pi_{\theta}||\pi_{\theta_{0}})$ to denote $\mathbb{E}_{x}[\mathrm{KL}(\pi_{\theta}(\cdot|x)||\pi_{\theta_{0}}(\cdot|x))]$ for short. We compare vanilla model averaging methods with the reward penalty by considering different KL penalties in $\{0.05,0.1,0.2\}$. The results are shown in Figure 8. We can see that while a larger KL penalty can partially mitigate the forgetting issue, the model averaging is much more effective than the reward penalty in terms of the alignment-forgetting trade-off.
Figure 8: Comparison of model averaging with reward penalty for PPO.
C.3 Consistency of different combination ratios among various tasks
We try three patterns experiment given a base $\alpha\in\{0.2,0.3,0.4\}$ : (a) $\alpha_{1}=\alpha_{2}=\alpha_{3}=\alpha$; (b) $\alpha_{1}=\alpha_{2}=\alpha$, $\alpha_{3}=\alpha-0.1$, and (c) $\alpha_{1}=\alpha$, $\alpha_{2}=\alpha_{3}=\alpha-0.1$. We use $(\alpha|\alpha|\alpha)$, $(\alpha|\alpha|\alpha-0.1)$ and $(\alpha|\alpha-0.1|\alpha-0.1)$ to denote these three patterns, respectively. These results confirm that certain ratio combinations exceed the trade-off curve of vanilla model averaging, as displayed in Figure 9. Notably, some combination ratios consistently outperform the equal ratio across various benchmarks. This affirms the potential to identify consistent combination ratios that demonstrate superior performance across a broad spectrum of benchmarks in terms of alignment-forgetting trade-off.
Figure 9: Evaluation of different combination ratios.
C.4 Results of $\alpha=0.2$
The following results show that when we chose $\alpha=0.2$, MA and HMA consistently alleviate the alignment tax without sacrificing any alignment performance.
Figure 10: Illustration of $\alpha=0.2$ on vanilla model averaging
Figure 11: Illustration of $\alpha=0.2$ on HMA
Appendix D Implementation Details